View  Edit  Attributes  History  Attach  Print  Search

Timeline

Timeline for the PPTA project

This page contains a list of the major events for the PPTA project. These include changes to the instrumentation of software that affect the observations (in blue), major publications and other useful dates.

  • 02/2003 Commissioning of WBC system. Limited configs and still instrumental problems.
  • 02/06/2003 Commencement of Dick Manchester's Federation Fellowship Project "Precision Pulsar Timing and its Applications"
  • 01/07/2003 George Hobbs transferred from Bolton Fellowship to Post-doc in FF project
  • 10/2003 Maryam Hobbs employed as part-time research assistant
  • 09/11/2003 Removal of MB receiver, installation and commissioning of 1050cm receiver, installation of H-OH receiver with upgraded preamps.
  • 10-12/2003 More WBC commissioning tests. Most problems solved
  • 11/2003 - 02/2004 Ryan Clement employed as Summer Student working on RFI mitigation algorithms
  • 06/02/2004 First PPTA (P456) observations at Parkes
  • 21/05/2004 Discovery of configuration dependent time delays in the WBC (RNM email 19/06/2004)
  • 07/06/2004 Tejinder Uppal commenced as software engineer working on RFI mitigation project
  • 01/07/2004 John Sarkissian employed as 50% part-time research assistant
  • 07/2004 Russell Edwards commenced post-doctoral appointment
  • 07/09/2004 MB Receiver reinstalled
  • 09/2004 WBC first bin - last bin problem fixed
  • 12/11/2004 New feed polarisation parameters introduced, FITS Version 1.24.
  • 03/2005 Start of official data collection for the PPTA project
  • 30/4/2005 Maryam Hobbs part-time term appointment completed
  • 05/2005 Publication of Kesteven et al. Adaptive filters revisited: Radio frequency interference mitigation in pulsar observations
  • 06/06/2005 Joris Verbiest starts PhD
  • 06/06/2005 Tejinder Uppal's term appointment completed
  • 06/2005 Publication of Jenet et al. Detecting the Stochastic Gravitational Wave Background Using Pulsar Timing
  • 20/06/2005 Commissioning of PDFB1 (Prototype DFB, 256 MHz bw)
  • 10/08/2005 Diana Londish commenced as part-time research assistant
  • 11/2005 Input levels to PDFB1 reduced to avoid saturation problems.
  • 01/01/2006 Implementation by JR of fix to PDFB1 configs to correct PTU problem diagnosed by RTE. Effect was sawtooth drift pattern of ~200ns in 5 min of TOAs.
  • 06/2006 Publication of Hobbs et al. TEMPO2, a new pulsar-timing package - I. An overview
  • 31/05/2006 PDFB1 linearity tests by RNM, JR, AB.
  • 21/06/2006 Amplifiers preceding digitiser inputs removed.
  • 07/08/2006 More linearity tests. LEVCAL operation implemented to reduce SPD levels to 5.0. STT_OFFS modified to allow for nchan-dependent propagation delays through digitiser (indicated by PROC_CMD in the HISTORY table being set to PSRDFB:2006-08-07_14:55). Attenuator settings moved to digstats table. Gal l,b, PA, FA, parang, tel coords added to subint table.
  • 09/08/2006 Diana Londish's term appointment completed
  • 03/10/2006 Moved DIG_STAT table to end of FITS file to avoid CORDAT not keeping up problem.
  • 1/11/2006 Lucyna Kedziora-Chudczer commenced as part-time research assistant
  • 11/2006 Publication of Edwards et a. TEMPO2, a new pulsar timing package - II. The timing model and precision estimates
  • 21/11/2006 Russell Edwards resigns to take up high school teaching in Victoria
  • 12/2006 Publication of Jenet et al. ''Upper Bounds on the Low-Frequency Stochastic Gravitational Wave Background from Pulsar Timing Obsevations: Current Limits and Future Prospects"
  • 16/12/2006 Changed from MULTI_1 to HOH - central frequency from 1369 MHz to 1433 MHz
  • 03/2007 Sarah Burke starts PhD
  • 05/03/2007 Daniel Yardley starts PhD
  • 16/03/2007 First light for PDFB2, max bandwidth = 512 MHz.
  • 16/03/2007 PDFB2 BE delays set to (incorrect) PDFB1 values, i.e., 32 chan: 1.2969 us; 128 chan: 3.5547 us; 512 chan: 12.2813 us; 2048 chan: 46.8829 us. For 256 and 1024 chan, no correction was applied. Actual delays are d1 + d2, where d1 = 14.00 us with an unpredictable fluctuation of up to 60 ns, and d2 is bandwidth and channel dependent as follows: 256 MHz, 512 chan: 13.91 us; 256 MHz, 1024 chan: 17.20 us; 256 MHz, 2048 chan: 43.15 us. The actual delays fluctuate around these values by 100 ns or so. Delays for other bw/chan combinations are currently not determined.
  • Between 17/04/2007 and 01/05/2007 Changed FITS HDRVER from 2.5D to 2.10 for VO compatibility. Parallel operation of PDFB1 and PDFB2.
  • 10/04/2007 David Smith commenced as part-time programmer
  • 18/04/2007 Jonathan Khoo commenced as part-time programmer
  • 05/2007 Use of Tempo2 predictor files with PDFB2
  • 06/05/2007 Linearity tests for PDFB2 by RNM and JV. LEVCAL setting of rms to 10.0
  • 17/05/2007 Changed from H-OH to Multibeam - central freq. from 1433 to 1369 MHz
  • Around 19/05/2007 Start running PDFB1 in alt mode. Occasional instances of alt-mode TCS giving incorrect sign of bw (e.g., +256 for MULTI). Spectrum remains inverted despite bw sign.
  • 05/06/2007 Commissioning of 1GHz PDFB2 system. The actual d2 delays (see 16 March above) are halved. For 1024 MHz, 1024 chan the actual delay is 4.59 +/- 0.1 us.
  • 06/2007 Implementation of pazi for pre-processing
  • 20/06/2007 Xiaopeng You completes PhD
  • 21/06/2007 Applied PDFB2 delays changed to the values appropriate for the 512 MHz system (see 5 June above) - i.e., twice the actual value!
  • 06/2007 Publication of You et al. Dispersion measure variations and their effect on precision pulsar timing'
  • Around 01/07/2007 Reported rms levels for PDFB2 inadvertently reduced by sqrt(8). Input power increased by about 10db to compensate, bringing rms levels back to 10.
  • 02/08/2007 Par files in ppta.db updated (V1.3)
  • 09/2007 Defaults for TZMJD etc. for T2 predictors
  • 09/2007 David Champion commenced post-doctoral appointment
  • 09/2007 Publication of You et al. An improved solar wind electron-density model for pulsar timing
  • 23/10/2007 Discovered reason for sudden change in DFB2 digitiser levels on 01/07/2007. RMS levels multiplied by sqrt(8) to restore to correct value. LEVCAL now setting input levels correctly.
  • 19/11/2007 PDFB2 configurations changed from gain 7 to gain 6; has effect of reducing uncalibrated amplitudes by a factor of 16.
  • 02/12/2007 John S. to use '-L' option to bin-zap automatically when pre-processing
  • 02/12/2007 WEW updates PDFB2 FITS version to 2.17. This includes BE_DELAY - the backend propagation delay from the digitiser input. STT_OFFS unchanged, i.e., still sum of delay to pulse phase zero and BE delay.
  • 17/12/2007 Pulse phase instabilities in PDFB1 data. PDFB2 appears OK. Possibly fixed by power cycle of whole PDFB1 system 20/12/07, 03UT.
  • 29/12/2007 Redefinition of STT_MJD by psrchive removed. Previously middle of 1st subint, now original value.
  • 01/01/2008 Start of GH's QEII Fellowship.
  • 16/01/2008 New PSRFITS definition including BE_DELAY
  • 20/02/2008 First light PDFB3 (V0)
  • 01/05/2008 PDFB3 schedule files in /psr1/tcs/sched.
  • 30/06/2008 David Smith's term appointment ended.
  • 29/07/2008 New PDFB3 boards installed by Andrew Brown (MJD 54676). Most recent firmware with PTU on-board. (V1)
  • 29/07/2008 New single-board system, PDFB4, commissioned.(V0)
  • 12/08/2008 At 0545UT loss of 5 MHz phase lock on PDFB3, leading to random changes in pulse phase. Largely fixed by JR inserting 8m cable in 5 MHz line.
  • 22/08/2008 Second PDFB3 board (A2/B2) replaced with spare unit.
  • 27/08/2008 Gain stability problem with B channel of PDFB3 and PDFB4 noticed (JR). Gain variations of ~1db on ~minute timescales.
  • 31/08/2008 PDFB1 decommissioned
  • 03/09/2008 B-channel gain stability problem shown to be in Eng Rack splitter amplifier (Mal Smith). Amplifier bypassed since approx unity gain overall.
  • 08/09/2008 Original PDFB3 second board restored since spare unit (installed 22/08/2008) had digitiser problem.
  • 10/09/2008 5 MHz phase-lock problem still intermittently present. Another 8m cable added by JR.
  • 22/09/2008 PDFB4 software and firmware updated (AB). Reset bug introduced so that folding period not updated leading to phase drift during observation. Also intermittent phase jumps introduced due to sync problem in PTU firmware. (V1)
  • 29/09/2008 PDFB4 hardware and firmware changed to fix PTU sync problem.
  • 12/10/2008 PDFB4 reset bug fixed ~0500UT 54751.30 (AB) (V2, First reliable PDFB4 data)
  • 16/10/2008 PDFB3 software upgraded to be same as PDFB4 (WEW) (V2)
  • 31/01/2009 Joris Verbiest submits PhD
  • 01/01/2009 Following the leap second at 0000UT on 01/01/09 and until 2350UT on 03/01/09 for PDFB3 and until 0000UT on 05/01/09 for PDFB4, recorded start times are late by 1 second, i.e. 1 second should be subtracted from stt_smjd for these files.
  • 11/02/2009 UT0600: PDFB4 control software upgraded to use FITS def V3.9.
  • 17/02/2009 Stefan Oslowski starts PhD
  • 26/04/2009 UT2300 PDFB3 delay compensation activated. (V4) MJD 54947.96 [MTK: NB this means correction for two-bin delay; configuration dependent delays still present.]
  • 12/06/2009 Lucyna Kedziora-Chudczer's term appointment ended.
  • 27/07/2009 10/50cm receiver reinstalled with 50cm band at 700-764 MHz, reconditioned 10cm preamps.
  • 29/07/2009 Entry for J0437-4715 in ppta.db updated to correct drift in PDFB pulse phase during obs
  • 31/08/2009 From this date, occasional incorrect DM (= 1) in PDFB3 & PDFB4 file headers. Data are OK.
  • 31/08/2009 David Champion left to take position at MPIfR, Bonn.
  • 02/09/2009 New version of PDFB3 control software installed. Uses Ver 4.0 of PSRFITS and other changes including to level control system. PSREPHEM table replaced by PSRPARAM (block text) table. (V5)
  • 02/09/2009 APSR/PDFB3 frequency leakage problem fixed by Andrew Brown. Start of good APSR data.
  • 15/09/2009 Ankur Chaudhary commenced as part-time research assistant
  • 28/10/2009 Eng Rack distribution amplifier replaced with three-way splitter. 6db pad on inputs maintains levels close to previous values.
  • 18/11/2009 Changed default predictor to tempo2 for psrchive at Epping
  • 14-24/12/09 Pks clock problems. Clock lost lock at MJD 55179.50 (Dec 14), lock restored at MJD 55181.206 (Dec 16) at +46.6 us. Reset to -0.97 us at MJD 55189.173 (Dec 24).
  • 07/01/10 DM=1 problem which has existed for PDFB3,4 data since 21/09/2009 fixed ~UT0200. Only pulsars with DM1 entries in par file were affected. pkccc3 and pkccc4 updated (~UT1000) to take DM1 into account.
  • 09/02/10 J2241-5236 added to PPTA schedule
  • 03/05/10 PDFB3 and PDFB4 firmware updated at UT0415 (MJD 55319.18) to remove delays, i.e. to make time-stamp refer to time of signal arrival at digitiser input. In case of PDFB4, the PTU was also changed to remove the two-bin delay. (V6,V3)
  • 30/06/10 CPSR2 and PDFB2 decommissioned.
  • 13-15/08/10 PDFB4 moved to new rack
  • 16/8/10 New firmware for PDFB3,4 (V7,V4) at 11UT, 54755.46
  • 17/8/10 PDFB3,4 firmware restored to V6,V3 at 06UT, 54756.25. J2241-5236 added to ppta.db and schedule files.
  • 07/09/10 J1017-7156 added to ppta.db and schedule files.
  • 18/9/10 10cm files with RF 3094 MHz and 20cm files with wrong sign for bandwidth identified by JER (JER and RNM emails of 2/8/10). These files fixed on Epping disks using:
psredit -m -c ext:obsfreq=3094.0; pam -m -o 3094.5
psredit -m -c bw=-256; pam -m -o 1368.875
(See RNM email of 14/9/10 for details)
  • 13/10/10 Switch matrix for IF control installed. CASPSR installed.
  • 18/10/10 Ryan Shannon joined project as postdoc.
  • 01/01/11 J1732-5049 dropped from regular PPTA observing
  • 08/2/11 Tempo2 updated at Parkes after a bug was found in the use of predictors (the site velocity was sometimes set to infinity when using a tempo2 ephemeris)
  • 25/2/11 0440UT (MJD 55617.2) PDFB3,4 updated control software in pkccc3,4 (WEW). PSRFITS V4.9
  • 26/2/11 ~06UT PDFB3 reverted to previous version since problems with APSR level control.
  • 27/2/11 Remaining files affected by 3094 MHz problem fixed on Epping disk
  • 20/3/11 PDFB3 updated control software reinstalled. TCS updated to include CASPSR.
  • 24/3/11 (MJD 55644) New ephemerides in ppta.db for all 22 pulsars.
  • 25/3/11 10cm data bad from UT0930 to UT1730 due to switch matrix problems related to CASPSR.
  • 1-24/3/11 Major shutdown. Snake cables on MB receiver replaced. Problems at cable-connector interface, causing intermittent bad connections.
  • 28/3/11 The A and B cables for the MB receiver were switched in the focus cabin during the shutdown. This was detected and repaired on 28/3/11. Data from MJD 55644.0 to 55648.2 are affected.
  • 6-10/6/11 IPTA Student Workshop in Morgantown WVa. RNM and RMS present.
  • 12-17/6/11 IPTA Science Meeting at Snowshoe Resort WVa. RNM and RMS present. GH gave talk via skype.
  • 29/10/11 PPTA logo design from Mike Keith adopted
  • 18/11/11 New par file for J1017-7156 from MJKeith
  • 18/11/11 10cm LO error from UT 06:00 - 17:05 (MJD 55883.25 - 55883.712). Effective RF 3094 MHz.
  • 01/12/11 Occasional loss of 5 MHz phase lock on PDFB4. Cable length adjusted by BP on 20/12/11.
  • 16/12/11 New versions of pdfb3 and pdfb4 (latest) with V5.0 of PSRFITS installed ~UT 09:30.
  • 6/4/12 5 MHz phase lock problems on PDFB4.
  • 3/5/12 Recurrence of loss of 5 MHz phase lock on PDFB4. Cable length adjusted by Tim Ruckley and John Sarkissian on 3/5/11 (20 metres of cable added)
  • 16/5/12 Brett P. removed 20m of extra cable on PDFB4 5 MHz line (5m left in). Fibre media adaptors in BAT line to PDFB4 removed and replaced by direct copper connection.
  • 23/5/12 Recurrence of loss of 5 MHz phase lock on PDFB4. Cable length adjusted by John Sarkissian and George Hobbs (5m removed and 20m added) at ~0400 UT.
  • 30/5/12 pdfb3 and pdfb4 cccversion "latest" updated to fix search-mode 8-bit 4-pol offset.
  • 14/6/12 PDFB4 removed and sent to Epping to fix connector problem caused by BAT line changes. Reinstalled 3 July.
  • 18/6/12 to 22/6/12 IPTA Winter School in Sydney. Organised by PPTA collaboration members. Highlight of school was Pulse@Parkes timing masterclass.
  • 25/6/12 to 28/6/12 IPTA Science Meeting in Kiama. Organised by PPTA collaboration members. Talks by Bailes, Burke Spolaor, Coles, Keith, Manchester, Oslowski, Ravi, Shannon, van Straten.
  • 14/8/12 Maser relocated. GPS system OK, but CommonView failed. Also changes to BAT wiring after clock; effect should be 20ns or less.
  • 17/8/12 CommonView GPS transferred to old single-frequency GPS.
  • 12/9/12 CommonView GPS restored to dual-frequency system.
  • 25/3/13 Power failure in tower etc UT 15:45 (MJD 56376). Among other things, clock maser failed. Systems restored by John S and Brett P ~18:00UT. ~600ns offset in Pks clock. Reset to ~20ns on 26/3/13 at ~22h UT.
  • 12/4/13 Mike Keith and George Hobbs updated the catalogue for APSR and CASPSR (noting that J0437-4715 incorrectly had an old A1DOT in the previous ephemeris)
  • 9/5/13 Maser glitch ~6us at UT 2240 (MJD 56421.93). Reset by Brett P. at approx UT 0245 10/5/13 (MJD 56422.12)
  • 14/5/13 New version of pdfb3 and pdfb4 installed (cccversion latest). Fixes problems with search mode: spurious blocks at EoF for 1-bit mode, first sample zero, channel offset for first block. Now uses PSRFITS V5.1 with nbit=1,2,4,8 all written in B mode.
  • 17-21/6/13 IPTA Student Workshop in Krabi. Ryan Shannon attended and gave tutorial, Jingbo Wang and Xingjiang Zhu attended. Dick Manchester attended Friday.
  • 24-28/6/13 IPTA Science Meeting in Krabi. Attended by George Hobbs (talk), Dick Manchester (talks), Ryan Shannon (talk), Linqing Wen, Jingbo Wang and Xingjiang Zhu. Mike Keith presented a talk via proxy GH.
  • 7/7/13 Xinping Deng and Jingbo Wang both graduated with PhDs.
  • 14/08/13 Updated pks2gps.clk at Epping (this was very out of date - last update in Oct 2012)
  • 01/09/13 A few days up until this date we were updating the observing software to remove dependency on /psr1. This required updates to psrcat, tempo2 and T2psrcoeff. Various issues were found including 1) pulsars with P0 and P1 instead of F0 and F1 were incorrectly converted for the ephemeris and 2) a bug was introduced into T2psrcoeff so that it did not include the PREDLEN parameter (fixed on the evening of 1/9/13). The various catalogue files were all moved to a new location and the most up to date versions of tempo2 etc. installed.
  • 28/09/13 Matthew Kerr joins the project
  • 14/01/14 Recognised that P456_50cm_DFB3.sch had LEVCAL mode for post-obs cals for J0437, J1744 and J1909. Should be CAL mode so atten not reset. Schedule fixed.
  • 14/01/14 Updated psrcat so that the ephemeris does not contain "tabs" (only spaces)
  • 21/01/14 New versions of pdfb3 and pdfb4 installed. New version is labelled "current". Previous version "latest" now labelled "previous".
  • 26/02/14 During the early hours of the morning, the internal power supply for the maser failed. Brett has fixed it, but there will be a step-change which will be measured
  • 27/02/14 Temporary replacement supply. Clock reset about 3 us off, but still drifting slowly.
  • 17/03/14 "current" versions of pdfb3/4 replaced to correct spurious stt_date and stt_time entries at end of main header. Old "current" versions changed to "previous"
  • 05/14 to present Major problems with pdfb3. pdfb3 could not sync both FPGA boards making pulsar timing modes inoperable. Search mode (and spectral-line) modes continue to function. PPTA observations continue with CASPSR and PDFB4 only.
  • 06/14 2014 IPTA meeting held in Banff, Canada. Talks given by Matthew Kerr, George Hobbs, Daniel Reardon, Caterina Tiburzi, Vikram Ravi, Dick Manchester, Ryan Shannon, XIngjiang Zhu, Jingbo Wang, Shi Dai, and Bill Coles. Linqing Wen also attended.
  • 15/08/14 "current" versions of pdfb3/4 correlator programmes updated to have CTR_FRQ definition in history table conform to psrfits definition
  • 25/10/14 "current" at commencement of 20cm session on 25 Oct spectral kurtosis turned off on CASPSR.
  • 16/01/15 Updated the tempo2 files (EOPC, pks2gps and gps2utc) at Parkes. They were many years out of date.
  • 10/02/15 RMS observed in experimental 50cm modes with variable centre frequency (that may not agree with headers in fits files). Data should be considered unreliable
  • 13/04/15 WvS updated 3100+732.cmd to better configure the attenuators for the power levels required by CASPSR
  • 28/02/15 Alectown Optus 4G station powered -- causing serious problems to 50 cm observations.
  • 20/07/15 RMS updated DR1 processing to zap 2670-2690 MHz apparent 4G RFI.
  • 06/07/16 DFB4 moved to DFB3 rack. Delay changes expected. Bandpass observed to be different before and after move. Delay measurements include observation of J0437-4715 at 10cm with DFB4 and CASPSR (unchanged) before and after, as well as modulated cal scheme (multiple DFB4 configurations, pre-and-post; last configuration before/after pdfb4_2048_1024_1024) and by recording baseband data with CASPSR with the DFB4 polA and polB placed into CASPSR polA.
  • 09/07/16 Translator problems resulted in session with focus disabled during 10/50 observations. Receiver position optimized for 45 degrees elevation.
  • '16/08/16 CASPSR moved to complete backend movement to accommodate SETI equipment. Move took place on 16-17. Pre and post move delay measurements taken on the 15th and 19th (see email from Jimi Green 05/10/16).
  • 13/09/16 Maser failed. Check electronic log for details of corrupted observations. A new maser was installed on 16/09/16.

05/12/16 Multibeam returns: Initial observations have MB signal through port 4 of the switch matrix versus port 3. Timing properties not predicted to be different.

17/01/17 CASPSR was found to not have had the leap second added at the start of the year. Data from 1st Jan until this date will therefore be affected.

25/07/17 The version of PSRCAT software updated on CASPSR

31/08/17 The version of PSRCAT software updated on joffrey (used by TCS and the DFBs)